Daily Report

CPUC/ANF Environmental Monitoring Activities

Pacific Pipeline Construction Project

Aspen Environmental Group Wednesday, December 02, 1998

The following is a brief summary of the environmental monitoring activities that were completed on December 02, 1998. This summary is based on reports directly received from the field and has not been edited. If you have any questions please contact (818) 597-3407. Please note the item number and the correction that should be made. This report is intended to serve as a communication document and as an instrument for quality control.

Construction Highlights

12/02/98 (Wednesday)

(EMP 15.35). Welding, x-ray and coating continued from EMP's 15.90 to 15.71. Electrical work continued at Grapevine Pump Station. On Tejon Bench backfill continued to 16.89 and excavation continued working north from EMP 18.09. On Lebec Rd., excavation through the road continued working south from EMP 21.53. The hydrostatic testing crew welded together an 8" pipe and began to transfer water into the segment to be tested from EMP 25.43 to 21.75. Ron Vance, with the County Roads Dept. visited the site where the I-5 bore at EMP 16.49 will begin soon and informed the contractor that the dirt road that extends down to the site from the freeway is absolutely not to be used.

Segment 2B (Castaic to Sylmar): Report for 12-2-98. At Feedmill Rd and the Santa Clara river drilling crews have pulled the 1200' long 6" wide sleeve for the fiber optic conduit. Further, crews will begin a general clean up of the area including removing berms along the river bank. Fiber optic cable crews were pulling string through the duct between Pico Cyn and Calgrove Blvd. Trenching crews are nearing completion of the ditch and should complete work by Friday. Pipe crews have material in the ditch to EMP 71.8 with welders, coating and x-ray crews working nearby. Three hundred feet of fiber optic duct and slurry were completed to EMP 72.3 while asphalt was completed to EMP 72.4. Crews removed steel plates and continued hauling away spoils from the right of way.

Segment 3A & 3C (Sylmar to Los Angeles): Report for 2 December 1998: 3A: Fiber optic crews pulled cable between Grandview and Bekins Way today. Backfill crews spent the day working at Tujunga Wash. This included padding, and the fiber optic conduit installation on the bridge. The bore at Tuxford was slowed because the augers got stuck in the casing pipe. Crews plan on digging to the casing from the exit site. Welding, x-ray, and coating continued work at Pacoima Wash. Ditch crews dug more trench for the tie in to that bridge. 3C: Crews began pushing pipe in to the casing pipe under Hubbard Avenue today. The final slurry was poured in the ditch between Balboa Boulevard and the aqueduct crossing.

Inspection Summary

In the past 24 hours, 3 inspection event(s) were reported for 3 spread(s). Of these inspections, 0 were non-compliance event(s). Also 0 non-compliance resolution report(s) was resolved.

Non-Compliance Reports (NCR's) Issued

No new non-compliance reports were issued during the last reporting period.

Non-Compliance Resolution Reports (NCRR's) Issued

No new non-compliance reports were issued during the last reporting period.

Summary of Follow-up Items for Today

No follow-up items were reported for today.

Back to Daily Reports

Back to MMCRP Home Page